OpenStudy (henrietepurina):

PLEASE HELP ON ALGEBRA 1 QUESTION Don't just give answer.. explain plz :) Medal + Fan

OpenStudy (henrietepurina):

3x3y + 12xy - 3x2y - 12y Part A: Rewrite the expression so that the GCF is factored completely.

OpenStudy (asnaseer):

Ok - look carefully at all the terms in your equation - can you see anything that appears in ALL the terms?

OpenStudy (asnaseer):

e.g. does \(x\) appear in every term in your expression?

OpenStudy (henrietepurina):

Oh gimme a sec, the equation looks a bit messed up :D

OpenStudy (henrietepurina):

It's really supposed to be (3x^3)y + 12xy - (3x^2)y - 12y

OpenStudy (asnaseer):

Is this your equation?\[3x^3y+12xy-3x^2y-12y\]

OpenStudy (henrietepurina):

Yeah :)

OpenStudy (asnaseer):

Ok - so lets start with \(x\) - does it appear in EVERY term of the expression?

OpenStudy (henrietepurina):

No, not in the last one

OpenStudy (asnaseer):

good - so that rules it out as a GCF

OpenStudy (asnaseer):

now look at \(y\) - does it appear in EVERY term of the expression?

OpenStudy (henrietepurina):

Ah.... YES

OpenStudy (asnaseer):

so now what you need to do is to factor out a \(y\) from each term - do you know how to do that?

OpenStudy (henrietepurina):

No, can you teach me :D

OpenStudy (texaschic101):

now this is exactly how your supposed to help someone with their homework...good job

OpenStudy (texaschic101):

don't forget the common number...not just the variable

OpenStudy (henrietepurina):

Are you complementing me or asnaseer ;P

OpenStudy (asnaseer):

@texaschic101 - thx :) Ok - first lets rewrite each term as follows:\[3x^3y+12xy-3x^2y-12y=y(3x^3)+y(12x)-y(3x^2)-y(12)\]does that make sense so far?

OpenStudy (texaschic101):

I am complimenting both of you

OpenStudy (henrietepurina):

Thanks :) And, ah... yes it does, you just removed the y's :D

OpenStudy (asnaseer):

Good - now we can take this COMMON \(y\) write out and rewrite the expression as follows:\[3x^3y+12xy-3x^2y-12y=y(3x^3)+y(12x)-y(3x^2)-y(12)\]\[=y(3x^3+12x-3x^2-12)\]agreed?

OpenStudy (henrietepurina):

Did you take the equation without the y's and multiplied it all by a single y?

OpenStudy (asnaseer):

yes - the \(y\) outside will multiply by each term inside the brackets

OpenStudy (asnaseer):

think of it like this:\[2*3=2*(1+2)=2(1+2)=2*1+2*2=2+4=6\]

OpenStudy (henrietepurina):

Ah ok....

OpenStudy (asnaseer):

this is called distributing the factor

OpenStudy (asnaseer):

Ok - now look at what remains inside the brackets:\[3x^3+12x-3x^2-12\]Can you see if there is any number that divides into every single term?

OpenStudy (henrietepurina):

Yes 3

OpenStudy (asnaseer):

good - so now we do the same trick and write this as:\[3x^3+12x-3x^2-12=3(x^3)+3(4x)-3(x^2)-3(4)\]\[=3(x^3+4x-x^2-4)\]agreed?

OpenStudy (henrietepurina):

Yes :)

OpenStudy (asnaseer):

great - no lets combine everything that we done to get:\[3x^3y+12xy-3x^2y-12y=y(3x^3+12x-3x^2-12)=3y(x^3+4x-x^2-4)\]

OpenStudy (asnaseer):

I hope that all makes sense

OpenStudy (asnaseer):

the final term is supposed to be:\[=3y(x^3+4x-x^2-4)\]

OpenStudy (henrietepurina):

OH GOLLY SO MANY DIGITS, but uh yeah it does make sense :D

OpenStudy (asnaseer):

good :)

OpenStudy (henrietepurina):

Alrighty'o, so overall the rewritten the expression so that the GCF is factored completely is 3y(x^3+4x−x^2−4)? @asnaseer

OpenStudy (henrietepurina):

(refresh page to remove question marks)

OpenStudy (asnaseer):

yes :)

OpenStudy (henrietepurina):

Thanks, but do you mind helping me on another one?

OpenStudy (asnaseer):

yw :) <<<---- please post each question separately on the left I will try to help but if I am busy with other things then be assured that there are plenty of very good helpers on this site who will be more than happy to help :)

OpenStudy (henrietepurina):

Oh ok, will do :) Thanks a ton for the help :)

OpenStudy (asnaseer):

again - yw :)